Divide 124 into 4 parts which is in progression such that the product of the 1st and the 4th part is 128 less than the product of the 2md and the 3rd part

Grade:10

1 Answers

Arun
6 years ago
Dear students
Let the 4 parts be
a- 3d, a -d, a+d, a+ 3d
Sumnof all of this is 124
Hence
4a = 124
a = 31
(a -3d) (a+3d) = (a-d)(a+d) -128
a² - 9d² = a² - d² -128
8d² = 128
d² = 128/8 = 16
d = +4 or -4
Hence if d = +4
Then numbers are
19, 27, 35, 43
 
If we take d = -4, we get aame numbers but in reverse order.
 
Regards
Arun (askIITians forum expert)